# Binary superfluids: low-energy properties and dissipative processes from spontaneous emission of massive phonons

### v1

@@ Abstract
We discuss the low-energy properties of binary superfluids with density-dependent interactions. Adding an intra-species coupling that induces an explicit soft symmetry-breaking, we determine the background pressure and we show that the low-energy spectrum consists of a massless Nambu-Goldstone boson and a massive (pseudo) Nambu-Goldstone boson. When the background velocities of the two superfluids are transonic, the system is characterized by two distinct acoustic horizons: the hydrodynamic analogue of the black hole event horizon. The Hawking-like emission occurring at these horizons produces an effective friction on the fluids. We compute the viscosity-to-entropy ratios close to the two acoustic horizons, finding that the emission of pseudo Nambu-Goldstone bosons violates the bound conjectured by Kovtun, Son and Starinet.
@@ I Introduction
The current generation of quantum technologies opens up the possibility of investigating the physics of ultracompact objects, like neutron stars or black-holes Barcelo et al. (2005); Altman et al. (2021); Warszawski and Melatos (2011); Warszawski et al. (2012); Faccio et al. (2013); Poli et al. (2023); Bland et al. (2024); Yago Malo et al. (2024); Cipriani et al. (2024); Coviello et al. (2024); Magierski et al. (2024); Tu and Li (2024); Schützhold (2025), in analogue experiments involving quantum fluids with tunable degree of correlations. Concerning black holes, key aspects are the dissipative effects close to horizons, as characterized by the shear viscosity-to-entropy ratio $\eta/s$ Liberati et al. (2006a); Adams et al. (2012). This quantity is known to depend on the degree of correlations in the fluid and reaches a minimum in correspondence of a critical point Adams et al. (2012). Interestingly enough, this minimum is of the order of $\hbar/k_{B}$ in strongly correlated quantum fluids, $k_{B}$ being the Boltzmann constant, with a value that is similar in quite diverse systems, as for instance the quark-gluon plasma and ultracold Fermi gases, irrespective of their so different microscopic details. The lower bound $\eta/s=1/4\pi$ in units of $\hbar/k_{B}$, known from Kovtun, Son, Starinet (KSS) Kovtun et al. (2005), has been derived in different frameworks, i.e. gravity-dual theories Cremonini (2011), analogue gravity Barcelo et al. (2005), kinetic theories Mannarelli and Manuel (2010); L. Chiofalo et al. (2024), and more generally from Kubo relations Adams et al. (2012); Barcelo et al. (2005); Cremonini (2011). While the KSS bound is seen to hold for a remarkably large class of theories, specific situations show that it is non-universal Adams et al. (2012); Cherman et al. (2008). These include translationally invariant non-Fermi liquids in two dimensions Ge et al. (2020), the presence of higher curvature corrections in the dual gravitational theory and massive gravity in a higher derivative theory, such as the Gauss–Bonnet theory Sadeghi and Parvizi (2016). The latter introduces a lower bound $\eta/s_{GB}=(1/4\pi)(16/25)$, which in fact avoids violations of causality and of energy positivity Adams et al. (2012). Overall, lower bounds $\eta/s$ establish that the low-shear viscosity systems, i.e. nearly perfect fluids, are also strongly correlated, bearing important information to fully exploit the analogy between black hole horizons and quantum fluids. Additionally, whether only one universal $\eta/s$ bound does exist, comprising all possible theories, remains a relevant open question. It is therefore especially interesting to design a theoretical framework, amenable to realistic experimental implementations, to explore the degree of KSS-bound violation under the tuning of specific external parameters.
@@
Multicomponent quantum gases offer a versatile platform for investigating a wide range of phenomena, for instance due to the possibility to manage multiple levels simultaneously, and to the high tunability of both intraspecies and interspecies interactions. These features have been shown to allow access to novel phases of matter Cornell and Wieman (2002); Recati and Stringari (2022); Baroni et al. (2024), possibly supporting nontrivial topological defects Son and Stephanov (2002a); Kevrekidis et al. (2004); Bakkali-Hassani et al. (2021); Romero-Ros et al. (2024); Hamner et al. (2011); Farolfi et al. (2020); Richaud et al. (2023), as well as self-bound quantum droplets Petrov (2015); Semeghini et al. (2018). The first experimental realization of two overlapping quantum bosonic mixtures was achieved by sympathetic cooling of two integer spin states of 87Rb atoms Myatt et al. (1997). Subsequently, binary superfluids of fermions, of bosons as well as of a mixture of fermionic and bosonic gases have been explored to investigate various relevant phenomena for elementary physics Tuoriniemi et al. (2002); Rysti et al. (2012); Lewenstein et al. (2012); Inguscio and Fallani (2013); Ferrier-Barbut et al. (2014); Lepori et al. (2015); Lepori and Mannarelli (2019). Among them, we mention the Andreev-Bashkin effect Nespolo et al. (2017), analog black-holes  Fischer and Schutzhold (2004); Visser and Weinfurtner (2005); Liberati et al. (2006b); Berti et al. (2024) and bubble formation in first-order phase transitions Zenesini et al. (2024). The possible phases of multicomponent quantum gases are also of relevance to characterize the interior of compact stars Yago Malo et al. (2024). Inside them, superfluid neutrons should coexist with superconducting protons, moreover, if deconfined quark matter is present, baryonic superfluidity may be realized along with color superconductivity Shapiro and Teukolsky (1983); Alford et al. (1999); Bedaque and Schafer (2002); Kaplan and Reddy (2002); Alford et al. (2008); Anglani et al. (2014).
@@
In this work, we focus on a binary mixture of weakly interacting bosons at vanishing temperature. In each component, superfluidity arises when the global symmetry associated with the conservation of number of particles is spontaneously broken Annett (2004); Pethick and Smith (2008), resulting in the presence of massless Goldstone bosons in the low-energy spectrum, i.e. the phonons. When an intra-species Rabi coupling is added, such symmetry is explicitly broken Son and Stephanov (2002a); Lepori et al. (2018) and the low-energy spectrum of the binary mixture is modified. If the Rabi coupling has a subleading effect, the low-energy spectrum consists of both a massless and a massive phonon Pu and Bigelow (1998a, b). These degrees of freedom may be interpreted as density waves Hall et al. (1998), and they can be described as massless and massive scalar particles propagating on top of an emergent acoustic metric Unruh (1981); Bilic (1999); Visser and Molina-Paris (2010); Brout et al. (1995); Barcelo et al. (2005); Volovik (2003). This system is particularly interesting when the flow of one or both species is transonic, providing the hydrodynamic analog of black-holes Fischer and Schutzhold (2004); Liberati et al. (2006a); Berti et al. (2024). Similarly to astrophysical black-holes, Hawking-like emission of phonons occurs at the acoustic horizon, imprinting a signature in the density-density correlation functions Carusotto et al. (2008); Steinhauer (2016); Berti et al. (2024). This spontaneous phonon emission is an irreversible process that occurs at the expense of the fluid kinetic energy, and can be described as effective shear and bulk viscosities of the superfluid L. Chiofalo et al. (2024).
@@
Within this framework, we compute both the shear-to-entropy density ratio and the bulk-to-entropy density ratio associated to the spontaneous emission at the acoustic horizon of binary transonic superfluid, extending the previous results for a single species L. Chiofalo et al. (2024). We show how the saturation of the bound occurs for the spontaneous emission of massless phonons, while the bound is violated when a massive phonon is emitted. In fact, in the latter condition, the corresponding pressure and energy density decrease, reducing the effective shear and bulk viscosities with respect to the massless case. These findings are shown in the nonrelativistic limit and are directly applicable to homonuclear bosonic mixtures of alkali atoms with coupled hyperfine states Myatt et al. (1997); Semeghini et al. (2018).
@@
This paper is organized as follows. In Sec. II, we briefly review the analogue gravity set-up with a single-species boson gas. In Sec. III, we describe the general features of binary bosonic systems, while in Sec. IV we determine the ground-state pressure. The low-energy field theory of binary superfluids is developed in Sec. V. The Hawking-like emission of massive phonons is studied in Sec. VI and in Sec. VII we evaluate the resulting effective viscosity coefficients. We draw our conclusions in Sec. VIII. Two distinct and general derivations of the effective low-energy Lagrangian are given in the Appendix A.
Along all the work, we use natural units $\hbar=c=k_{B}=1$.
@@ II.1 General considerations
We consider a complex scalar field, $\Phi$, with non-vanishing chemical potential and local interactions that do not explicitly break the $U(1)$ global symmetry associated to the conserved number of particles. We assume that the bulk temperature of the system is negligible. For definiteness, we examine a simple model with contact self-interaction; for a more general system see the Appendix A. The corresponding Lagrangian density, hereafter the Lagrangian, is given by
@@
is the covariant derivative, depending on the medium properties, see for instance Son (2002) for more details. We identify $\mu$ with the boson chemical potential, while $v_{\nu}=\gamma(1,\bm{v})$ is the fluid four-velocity: $\gamma$ is the Lorentz factor, such that $v^{\mu}v^{\nu}\eta_{\mu\nu}=1$, with $\eta_{\mu\nu}=\text{diag}(1,-1,-1,-1)$ the Minkowski metric. The non-relativistic limit of Eq. (1) is the Gross-Pitaevskii (Hartree-Fock) Lagrangian describing a diluted and weakly interacting boson gas at vanishing temperature Pethick and Smith (2008); Dalfovo et al. (1999). The excitation spectrum over the superfluid ground-state can be calculated linearizing the Gross-Pitaevskii Lagrangian, this approach being equivalent to the Bogoliubov formalism Pitaevskii and Stringari (2003); Recati and Stringari (2022).
@@
showing that the chemical potential gives two distinct contributions to the Lagrangian: an explicit Lorentz symmetry breaking, due to the term linear in the chemical potential, and the effective mass shift $\mu^{2}/\gamma^{2}$. When $\mu=\gamma\,m$ the effective mass of the scalar field vanishes, signaling a spontaneous symmetry breaking. For $\mu>\gamma m$, the $U(1)$ global symmetry is indeed spontaneously broken  [^1] This is what occurs in the framework of the Bogoliubov theory for superfluid bosons, see for instance Volovik (2009)., and the minimum of the potential
@@
The fact that the potential depends only on $|\Phi|$ and not on its phase, invites to use the Madelung representation
@@
where the two real scalar fields $\rho$ and $\theta$ are called the radial and phase fields, respectively. In the Madelung representation the Lagrangian in Eq. (II.1) reads
@@
is conserved. In the ground state, the phase is homogeneous, therefore
@@
corresponds to the solution of the stationary condition that minimizes the potential in Eq. (II.1). This is equivalent to maximize the ground state Lagrangian. Since the grand potential per unit volume is given by
@@
$P=\frac{(\mu^{2}-m^{2}\gamma^{2})^{2}}{4U\gamma^{4}}\,{\color[rgb]{0,0,0}.}$ (15)
Notice that this includes only the contribution of condensation and neglects temperature as well as quantum fluctuations. The velocity dependence, determined by the Lorentz factor, gives the relativistic analogue of the Bernoulli equation. Indeed, we have that
@@ II.2 Effective low-energy theory
where the dots indicate neglected cubic terms and higher, as well as surface terms. The term ${\cal L}_{0}(\rho,\partial_{\mu}\rho,\partial_{\mu}\theta)$ is the background pressure previously evaluated in Eq. (15). The linear term vanishes, because we are evaluating it at the stationary point of the action. The quadratic part of the Lagrangian reads
@@
and the leading order in $p/\tilde{m}$ consists in replacing
@@
is the adiabatic speed of sound. This quantity vanishes for $U\rho^{2}\to 0$, corresponding to the transition from the broken to the normal phase, while its upper bound is $\sqrt{1/3}$.
We observe that retaining the leading order in $p/\tilde{m}$ in Eq. (26) is equivalent to neglect the kinetic term $\partial_{\nu}\tilde{\rho}\partial^{\nu}\tilde{\rho}/2$ in Eq. (19), and treat $\tilde{\rho}$ as a homogeneous field. This will be useful in the development of the low-energy theory of binary superfluids. More general approaches to the effective low-energy theory are reported in the Appendix A. Now, we move to consider binary mixture of superfluids.
@@ III Binary boson superfluids
In the previous section, we have seen how it is possible to define an emergent acoustic metric for the long-wavelength excitations of a single superfluid. We now extend the same approach to a mixture composed of two flavors $A$ and $B$. In experiments with homonuclear gases, they may correspond to two atomic hyperfine states. In astrophysics, they may indicate the superfluid components in the neutron star interior. We take a scheme-independent approach and, as in the previous section, we employ a relativistic framework. We assume vanishing temperature and that the two fluids move with velocities $\bm{v}_{A}$ and $\bm{v}_{B}$ with respect to an external observer.
@@
III.1 Lagrangian of binary bosonic systems

### v2

@@ Abstract
We discuss the low-energy properties of binary superfluids with density-dependent interactions. Adding an intraspecies coupling that induces an explicit soft symmetry breaking, we determine the background pressure and we show that the low-energy spectrum consists of a massless Nambu-Goldstone boson and a massive (pseudo) Nambu-Goldstone boson. When the background velocities of the two superfluids are transonic, the system is characterized by two distinct acoustic horizons: the hydrodynamic analog of the black hole event horizon. The Hawking-like emission occurring at these horizons produces an effective friction on the fluids. We compute the viscosity-to-entropy ratios close to the two acoustic horizons, finding that the emission of pseudo Nambu-Goldstone bosons violates the bound conjectured by Kovtun, Son, and Starinet.
@@ I Introduction
The current generation of quantum technologies opens up the possibility of investigating the physics of ultracompact objects, like neutron stars or black holes Barcelo et al. (2005); Altman et al. (2021); Warszawski and Melatos (2011); Warszawski et al. (2012); Faccio et al. (2013); Poli et al. (2023); Bland et al. (2024); Yago Malo et al. (2024); Cipriani et al. (2024); Coviello et al. (2024); Magierski et al. (2024); Tu and Li (2024); Schützhold (2025), in analog experiments involving quantum fluids with a tunable degree of correlations. Concerning black holes, key aspects are the dissipative effects close to horizons, as characterized by the shear viscosity-to-entropy ratio $\eta/s$ Liberati et al. (2006a); Adams et al. (2012). This quantity is known to depend on the degree of correlations in the fluid and reaches a minimum in correspondence of a critical point Adams et al. (2012). Interestingly enough, this minimum is of the order of $\hbar/k_{B}$ in strongly correlated quantum fluids, $k_{B}$ being the Boltzmann constant, with a value that is similar in quite diverse systems, as for instance the quark-gluon plasma and ultracold Fermi gases, irrespective of their very different microscopic details. The lower bound $\eta/s=1/4\pi$ in units of $\hbar/k_{B}$, known from Kovtun, Son, and Starinet (KSS) Kovtun et al. (2005), has been derived in different frameworks, i.e. gravity-dual theories Cremonini (2011), analogue gravity Barcelo et al. (2005), kinetic theories Mannarelli and Manuel (2010); L. Chiofalo et al. (2024), and more generally from Kubo relations Adams et al. (2012); Barcelo et al. (2005); Cremonini (2011). While the KSS bound is seen to hold for a remarkably large class of theories, specific situations show that it is nonuniversal Adams et al. (2012); Cherman et al. (2008). These include translationally invariant non-Fermi liquids in two dimensions Ge et al. (2020), the presence of higher curvature corrections in the dual gravitational theory and massive gravity in a higher derivative theory, such as the Gauss–Bonnet theory Sadeghi and Parvizi (2016). The latter introduces a lower bound $\eta/s_{GB}=(1/4\pi)(16/25)$, which in fact avoids violations of causality and of energy positivity Adams et al. (2012). Overall, lower bounds $\eta/s$ establish that the low-shear viscosity systems, i.e. nearly perfect fluids, are also strongly correlated, bearing important information to fully exploit the analogy between black hole horizons and quantum fluids. Whether all theories predict that one universal $\eta/s$ bound exists, remains a relevant open question. It is therefore especially interesting to design a theoretical framework, amenable to realistic experimental implementations, to explore the degree of KSS-bound violation under the tuning of specific external parameters.
@@
Multicomponent quantum gases offer a versatile platform for investigating a wide range of phenomena, for instance due to the possibility to manage multiple levels simultaneously, and to the high tunability of both intraspecies and interspecies interactions. These features have been shown to allow access to novel phases of matter Cornell and Wieman (2002); Recati and Stringari (2022); Baroni et al. (2024), possibly supporting nontrivial topological defects Son and Stephanov (2002a); Kevrekidis et al. (2004); Bakkali-Hassani et al. (2021); Romero-Ros et al. (2024); Hamner et al. (2011); Farolfi et al. (2020); Richaud et al. (2023), as well as self-bound quantum droplets Petrov (2015); Semeghini et al. (2018). The first experimental realization of two overlapping quantum boson mixtures was achieved by sympathetic cooling of two integer spin states of 87Rb atoms Myatt et al. (1997). Subsequently, binary superfluids of fermions, of bosons as well as of a mixture of fermion and boson gases have been explored to investigate various relevant phenomena for elementary physics Tuoriniemi et al. (2002); Rysti et al. (2012); Lewenstein et al. (2012); Inguscio and Fallani (2013); Ferrier-Barbut et al. (2014); Lepori et al. (2015); Lepori and Mannarelli (2019). Among them, we mention the Andreev-Bashkin effect Nespolo et al. (2017), analog black holes and cosmology Fischer and Schutzhold (2004); Visser and Weinfurtner (2005); Liberati et al. (2006b); Berti et al. (2024), and bubble formation in first-order phase transitions Zenesini et al. (2024). The possible phases of multicomponent quantum gases are also of relevance to characterize the interior of compact stars Yago Malo et al. (2024). Inside them, superfluid neutrons should coexist with superconducting protons, moreover, if deconfined quark matter is present, baryonic superfluidity may be realized along with color superconductivity Shapiro and Teukolsky (1983); Alford et al. (1999); Bedaque and Schafer (2002); Kaplan and Reddy (2002); Alford et al. (2008); Anglani et al. (2014).
@@
In this work, we focus on a binary mixture of weakly interacting bosons at vanishing temperature. In each component, superfluidity arises when the global symmetry associated with the conservation of number of particles is spontaneously broken Annett (2004); Pethick and Smith (2008), resulting in the presence of massless Goldstone bosons in the low-energy spectrum, i.e. the phonons. When an intraspecies Rabi coupling is added, such symmetry is explicitly broken Son and Stephanov (2002a); Lepori et al. (2018) and the low-energy spectrum of the binary mixture is modified. If the Rabi coupling has a subleading effect, the low-energy spectrum consists of both a massless and a massive phonon Pu and Bigelow (1998a, b). These degrees of freedom may be interpreted as density waves Hall et al. (1998), and they can be described as massless and massive scalar particles propagating on top of an emergent acoustic metric Unruh (1981); Bilic (1999); Visser and Molina-Paris (2010); Brout et al. (1995); Barcelo et al. (2005); Volovik (2003). This system is particularly interesting when the flow of one or both species is transonic, providing the hydrodynamic analog of black holes Fischer and Schutzhold (2004); Liberati et al. (2006a); Berti et al. (2024). Similarly to astrophysical black holes, Hawking-like emission of phonons occurs at the acoustic horizon, imprinting a signature in the density-density correlation functions Carusotto et al. (2008); Steinhauer (2016); Berti et al. (2024). This spontaneous phonon emission is an irreversible process that occurs at the expense of the fluid kinetic energy, and can be described as effective shear and bulk viscosities of the superfluid L. Chiofalo et al. (2024).
@@
Within this framework, we compute both the shear-to-entropy density ratio and the bulk-to-entropy density ratio associated to the spontaneous emission at the acoustic horizon of binary transonic superfluid, extending the previous results for a single species L. Chiofalo et al. (2024). We show how the saturation of the bound occurs for the spontaneous emission of massless phonons, while the bound is violated when a massive phonon is emitted. In fact, in the latter condition, the corresponding pressure and energy density decrease, reducing the effective shear and bulk viscosities with respect to the massless case. These findings are shown in the nonrelativistic limit and are directly applicable to homonuclear boson mixtures of alkali atoms with coupled hyperfine states Myatt et al. (1997); Semeghini et al. (2018).
@@
This paper is organized as follows. In Sec. II, we briefly review the analogue gravity set-up with a single-species boson gas. In Sec. III, we describe the general features of binary boson systems, while in Sec. IV we determine the ground-state pressure. The low-energy field theory of binary superfluids is developed in Sec. V. The Hawking-like emission of massive phonons is studied in Sec. VI and in Sec. VII we evaluate the resulting effective viscosity coefficients. We draw our conclusions in Sec. VIII. Two distinct and general derivations of the effective low-energy Lagrangian are given in the Appendix A.
Unless otherwise stated, we use natural units $\hbar=c=k_{B}=1$.
@@ II.1 General considerations
We consider a complex scalar field, $\Phi$, with nonvanishing chemical potential and local interactions that do not explicitly break the $U(1)$ global symmetry associated to the conserved number of particles. We assume that the bulk temperature of the system is negligible. For definiteness, we examine a simple model with contact self-interaction; for a more general system see the Appendix A. The corresponding Lagrangian density, hereafter called the Lagrangian, is given by
@@
is the covariant derivative, depending on the medium properties, see for instance Son (2002) for more details. We identify $\mu$ with the boson chemical potential, while $v_{\nu}=\gamma(1,\bm{v})$ is the fluid four-velocity: $\gamma$ is the Lorentz factor, such that $v^{\mu}v^{\nu}\eta_{\mu\nu}=1$, with $\eta_{\mu\nu}=\text{diag}(1,-1,-1,-1)$ the Minkowski metric. The nonrelativistic limit of Eq. (1) is the Gross-Pitaevskii (Hartree-Fock) Lagrangian describing a diluted and weakly interacting boson gas at vanishing temperature Pethick and Smith (2008); Dalfovo et al. (1999). The excitation spectrum over the superfluid ground state can be calculated linearizing the Gross-Pitaevskii Lagrangian, this approach being equivalent to the Bogolyubov formalism Pitaevskii and Stringari (2003); Recati and Stringari (2022).
@@
showing that the chemical potential gives two distinct contributions to the Lagrangian: an explicit Lorentz symmetry breaking, due to the term linear in the chemical potential, and the effective mass shift $\mu^{2}/\gamma^{2}$. When $\mu=\gamma\,m$ the effective mass of the scalar field vanishes, signaling a spontaneous symmetry breaking. For $\mu>\gamma m$, the $U(1)$ global symmetry is indeed spontaneously broken, and the minimum of the potential
@@
The fact that the potential depends only on $|\Phi|$ and not on its phase, invites us to use the Madelung representation
@@
where the two real scalar fields $\rho$ and $\theta$ are called the radial and phase fields, respectively. In the Madelung representation the Lagrangian in Eq. (II.1) reads as
@@
is conserved. In the ground state, the phase is homogeneous; therefore
@@
corresponds to the solution of the stationary condition that minimizes the potential in Eq. (II.1). This is equivalent to maximizing the ground state Lagrangian. Since the grand potential per unit volume is given by
@@
$P=\frac{(\mu^{2}-m^{2}\gamma^{2})^{2}}{4U\gamma^{4}}\,.$ (15)
Notice that this includes only the contribution of condensation and neglects temperature as well as quantum fluctuations. The velocity dependence, determined by the Lorentz factor, gives the relativistic analog of the Bernoulli equation. Indeed, we have that
@@ II.2 Effective low-energy theory
where the dots indicate neglected cubic terms and higher, as well as surface terms. The term ${\cal L}_{0}(\rho,\partial_{\mu}\rho,\partial_{\mu}\theta)$ is the background pressure previously evaluated in Eq. (15). The linear term vanishes, because we are evaluating it at the stationary point of the action. The quadratic part of the Lagrangian reads as
@@
and the leading order in $p/\tilde{m}$ consists of replacing
@@
is the adiabatic speed of sound. This quantity vanishes for $U\rho^{2}\to 0$, corresponding to the transition from the broken to the normal phase, while its upper bound is $\sqrt{1/3}$. The obtained effective field theory is valid for momenta lower than $\tilde{m}$, and it is equivalent to the hydrodynamic description of a superfluid, holding at distances larger than the healing length Dalfovo et al. (1999)
$\xi=\frac{1}{\sqrt{2}mc_{s}}\,.$ (29)
By combining the nonrelativistic limit of the sound speed $c_{s}\simeq\sqrt{U\rho^{2}/(2m^{2})}$ with the Eq. (20), we have
$\xi\simeq\frac{\sqrt{2}}{\tilde{m}}\,,$ (30)
showing that distances larger than $\xi$ correspond to phonon momenta smaller than $\tilde{m}$. We observe that retaining the leading order in $p/\tilde{m}$ in Eq. (26) is equivalent to neglecting the kinetic term $\partial_{\nu}\tilde{\rho}\partial^{\nu}\tilde{\rho}/2$ in Eq. (19), and treat $\tilde{\rho}$ as a homogeneous field. This will be useful in the development of the low-energy theory of binary superfluids. More general approaches to the effective low-energy theory are reported in the Appendix A. Now, we move to consider binary mixture of superfluids.
@@ III Binary boson superfluids
In the previous section, we have seen how it is possible to define an emergent acoustic metric for the long-wavelength excitations of a single superfluid. Here, we extend the same approach to a mixture composed of two flavors $A$ and $B$. In experiments with homonuclear gases, they may correspond to two atomic hyperfine states. In astrophysics, they may indicate the superfluid components in the neutron star interior. We take a scheme-independent approach and, as in the previous section, we employ a relativistic framework. We assume vanishing temperature and that the two fluids move with velocities $\bm{v}_{A}$ and $\bm{v}_{B}$ with respect to an external observer.
@@
III.1 Lagrangian of binary boson systems